OpenStudy (anonymous):

find the sum of this series. \[\sum_{n=1}^{\infty} (-2/5)^n\]

ganeshie8 (ganeshie8):

start by finding the first term and common ratio

OpenStudy (anonymous):

hmmm... common ratio: -2/5 and first term: 1?

ganeshie8 (ganeshie8):

you're right about common ratio = -2/5 first term is wrong

ganeshie8 (ganeshie8):

for the first term, plugin n = 1 : \[a_n = \left(-\dfrac{2}{5}\right)^n\]

ganeshie8 (ganeshie8):

\[a_1 = \left(-\dfrac{2}{5}\right)^1 = -\dfrac{2}{5}\]

ganeshie8 (ganeshie8):

so both first term and common ratio are -2/5 still remember the infinite sum formula ?

ganeshie8 (ganeshie8):

a = -2/5 r = -2/5 plugin the values in infinite sum formula

OpenStudy (anonymous):

it's -2/7?

ganeshie8 (ganeshie8):

thats right! you may use wolfram to double check http://www.wolframalpha.com/input/?i=%5Csum_%7Bn%3D1%7D%5E%7B%5Cinfty%7D+%28-2%2F5%29%5En

OpenStudy (anonymous):

Thank you soo much! You're amazing! :)